Circular No. 5962

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

COMET SHOEMAKER-LEVY (1994d) On Mar. 23 C. S. Shoemaker reported her discovery of a comet on

Palomar 0.46-m Schmidt films taken by D. H. Levy, D. Williams and herself on Mar. 14. The images were significantly trailed on the 10-min exposures. The coma was slightly asymmetric toward the southeast, and there was a moderately strong condensation. After several unsuccessful attempts using search ephemerides by the undersigned, Shoemaker, Levy and H. E. Holt have recovered the comet, which is now only slightly condensed with a short tail to the northeast. Measurers B. A. Skiff (Mar. 14), J. Mueller (Apr. 3).

1994 UT  α_{2000}  δ_{2000}  m_{1}
Mar. 14.35173  10^{h}54^{m}10.^{s}67  − 8^{o}23′54′′.2  13.0
14.39496  10 53 50.21  − 8 19 01.2
Apr. 3.15121  8 22 55.71  +27 03 28.0  14.2
3.18055  8 22 44.90  +27 05 30.3

Parabolic orbital elements by the undersigned:

T = 1994 May 27.295 TT  ω = 57.^{o}296
Ω = 166.559 }2000.0
q = 1.16089 AU  i = 131.302
1994TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Apr. 2 8^{h}30.^{m}10 +25^{o}40.^{′}8 0.745 1.448 111.^{o}3 40.^{o}0 14.0
4 8 17.84 +28 00.5 0.782 1.430 106.1 42.2 14.0
6 8 06.48 +30 03.0 0.821 1.412 101.2 44.0 14.1
8 7 56.01 +31 50.0 0.862 1.395  96.7 45.5 14.1

SUPERNOVA 1994I IN NGC 5194 B. Schmidt and R. Kirshner, Center for Astrophysics, report that unre-

duced spectra taken by J. Peters on the Whipple Observatory 1.5-m reflector (+ FAST spectrograph) and by G. Bernstein with the Multiple-Mirror Telescope (Red Channel) during Apr. 3.37–3.52 UT suggest that this is a type II supernova in its earliest phases. Broad absorption features are seen at 425 nm (Hγ), 471 nm (Hβ), 570 nm (He I) and 620 nm (Hα). The features attributed to Hβ, He I and Hα show two absorption components. Also seen are strong, narrow, Na D lines at the recession velocity of M51 with an equivalent width of 0.27 nm. The object’s spectrum showed considerable evolution during the course of the observations, with the lines becoming more pronounced over time.

1994 April 3 Brian G. Marsden